OpenStudy (anonymous):

Please help with a Quadratic homework problem. Write the quadratic equation whose roots are 5 and - 4, and whose leading coefficient is 2. (Use the letter x to represent the variable). ____ = 0

OpenStudy (luigi0210):

Work backwards

OpenStudy (anonymous):

here i will get you started :) \[c(x - r)(x - r)\] the c is the coefficient while the r is the root plug in the numbers and tell me what you get :)

OpenStudy (anonymous):

\[c(x - r_{1})(x-r_{2})\] just to show that you have to use both roots

OpenStudy (anonymous):

I multiplied the coefficient 2 by both roots but is unsure of my answer: 2^2-8-40=0

OpenStudy (anonymous):

is that -8x?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

i got something different :O you mind if i show you step by step and see if you can find mine or yours error?

OpenStudy (anonymous):

Please show me the step by step!

OpenStudy (anonymous):

well you should have setup the problem like this : \[2(x-5)(x+4)\] did you setup correctly?

OpenStudy (anonymous):

Yes, that is how my equation looks.

OpenStudy (anonymous):

Actually, no, mine was 2(x-5) (x-4) this was the example in my math lab.

OpenStudy (anonymous):

they gave that to you as a example? :( but it calls for a -4?

OpenStudy (luigi0210):

That's kind of right in a sense

OpenStudy (anonymous):

hmmmm!

OpenStudy (anonymous):

@Luigi0210 can you explain :) because now im confused on it

OpenStudy (luigi0210):

But yea, it needs a -4.

OpenStudy (luigi0210):

x=5 x=-4 Think of it this way. Ignore the coefficient for now

OpenStudy (luigi0210):

Now work backwards. Subtract 5 from both sides and add 4 to both: x-5=5-5 ------------------ x+4=-4+4 Giving you: (x-5)(x+4)

OpenStudy (luigi0210):

*=0

OpenStudy (anonymous):

so is are we wrong or did she get the wrong example?

OpenStudy (luigi0210):

(x-5)(x+4)=0 Now distribute this out

OpenStudy (anonymous):

because if she used (x-5)(x-4) we know what her problem is already

OpenStudy (luigi0210):

The example must be wrong.

OpenStudy (anonymous):

The example shows different numbers and neither of them are negative.

OpenStudy (anonymous):

did you give us the wrong numbers ?

OpenStudy (luigi0210):

The answer should be: 2x^2-2x-40

OpenStudy (anonymous):

yea what i got :)

OpenStudy (anonymous):

Ok, I was close but no cigar! Thank you guys and Luidi here is your medal.

OpenStudy (luigi0210):

Luidi? >.>

OpenStudy (luigi0210):

You're Welcome.

OpenStudy (anonymous):

You guys are great! Thank you!